You are here: Foswiki>Commissioning Web>Com20080413 (16 Apr 2008, NormCushing)Edit Attach

Commissioning Notes 20080413 UT

Observer: J. Hill, R. Green, R. Zhelem, I. Foppiani in the control room and a gaggle of programmers in Tucson
Telescope Operator: D. Huerta

Summary

To quote RFG at 5:10 AM: "This was actually good." (In understated exhaustion he was referring to the first LBT F/15 Gregorian images ever taken!)

We had some problems with HBS early, but managed to beat them into submission. Read some of my explanations below, I haven't updated the Issuetraks yet. After some very fun fooling around pointing at the moon, we managed to put Regulus on the IRTC detector using a spiral search in telescope offset, even though that same offset drives PCS crazy. We are now about 5 stars in to making the initial IRTC collimation model and pointing model. Had some minor problems with irs, OSS and PCS, but generally things worked well. Best image was sub-arcsec in H!, so we can say that the secondary has the correct figure. Richard has EF running, but can't deal with short exposures of bright stars because there is too much speckle noise.

Misc. Tests in the late afternoon

00:10 We opened the left shutter door with enclosure facing east and telescope at zenith to get daytime sky illumination for the alignment telescope. Ross checked the RY value with the alignment scope, and verified the centering of the primary relative to the secondary using multiple reflections. The nominal centered position of the SX M2 Hexapod is X=+5.50mm Y=+3.50mm Z=0 RX=-75arcsec RY=-280 arcsec, and SX M1 HPs were at X=-1.753 Y=+1.739 Z=-1.565 RX=+41.4 RY=-40.2

Nighttime Ops

02:00 Finished opening the dome and starting the telescope

02:30 Stopped and Restarted PCS as it was unresponsive.

02:32 Slew to moon 08 01.1 +22 48 from USNO ephemeris page, then offset slightly to look at the bright half.

02:38 Moon is visible in IRTC, and it is vaguely in focus!!!!!!! Corresponding MAT image is a bit too saturated to be very useful.

02:42 Telescope stops tracking in elevation. It appears to be the low oil pressure problem even though the oil was quite cold (-11 degC). Had to stop and restart MCSPU since the elevation servo appears to have stalled. This means restarting PCS, MCS as well. Time lost: half hour. But it was fun drift scanning across the moon!

The biggest problem was that HBS was just at the Ta value of -0.2 degC so that the required number of pumps was fluctuating between 0.9 and 1.1. The PLC was updating Ta at ~1 Hz, but Pump 4 requires several seconds to go between 0 and full speed, so the average number of pumps we were getting was 0.7. This was only enough to keep us at 110 Bar. JMH temporarily fixed the problem by increasing the flow parameters by 30% for nodes 2-5 of the flow lookup table. Oil tank temperature was reported by the PLC to be +9 degC, while Ta was near 0 degC and the coolant return temperature was -5 degC so there seems to be alot of discrepancy in temperature.

03:37 Restart OSS to clear up a UMAC communication problem.

03:30 Continue adjusting coordinates of moon with preset GUI until the mountains at the terminator near the north limb appeared in IRTC. We made an initial shift of the Primary mirror to reduce coma, but there is still plenty of coma present.

04:15 MAT image of Limb @ terminator saved as lunalunaluna.jpeg from Celia_0046.fit

04:16 slew to Regulus, temporarily adjust RA/DEC coordinates of Regulus to put it at the desired position on the MAT.

04:32 Lost AZ and EL drives. HBS had gone into alarm and shut off. David checks WinCC computer and reports that compensation pressure was OK. so we aren't sure why it shut down. Lost 12 minutes.

04:56 HBS shuts down again. Lost 30 minutes. Beside the compensation pressure problem, the HBS is having a cooling problem. Tank temp and reported temperature at pads is around 12-13 degC and increasing, but glycol output temperature is -0.5. So either we have insufficient flow in the cooling pumps or we've got a hot return channel back to the pump inlets.

05:30 Preset to Saturn, but couldn't find Saturn on IRTC.

05:40 Back to the moon, a target we could hit. We had trouble moving repeatably from the preset GUI, so we switched to doing telescope RA/DEC offsets with LBTtools

06:20 kill and restart OSS again.

06:39 Limb At terminator is at 374, 344 on WF MAT. EL=30deg (but just off the edge of NFMAT)

06:40 Slew back to Regulus

06:55 Point #2 in pointing log is where we started our spiral search. This was Regulus on the spot in WFMAT that we identified above.

07:08 365,359 on WFMAT is where we found Regulus with a spiral search.

Moving M1 Y more negative to compensate coma with bright edge near swing arm.

Celebrate with a bag of Gocciole 75% cacao biscotti! (Thanks Baru.)

*The asterisk indicates a useful image/time for collimation and pointing data.

*002505 is Regulus 0.5 mm extrafocal (-0.5 mm) in 0.01 sec. Needs negative spherical. Next image is the intrafocal, but we weren't centered in focus. SX primary global offsets were X=-2.1 Y=-2.9 Z=~0 RX=0 RY=0 at EL=~38 deg SX total positions are in image header from M1 and M2

Dave T. didn't fix the decimal point in LBTtools pointto, but preset failed because GCS LEFT was not running. (two errors ??)

Start GCSL.

07:36 Preset to WT10_365 Starts out on WFMAT at 370,373. (using the LBC red pointing model)

Load irtc.20080413.ptmod which is at the moment a copy of irtc.20080401.ptmod. Touch it on the to-station so David can browse, but editting it on lbtmu102 where PCS reads it. (Note this important point when updating pointing models.)

07:54 WT10_365 Starts out on WFMAT at 367,371 at EL=64.8

Adjusted IE -10 arcsec to -77 ===> 367,372

IE=-10 ===> 367,363

IE=+30: ===> 370,358 accidental repeat 367,357

08:06 IE=+25 IA +28 to -200: 365,357

08:09 IE=+20 IA=-140: 362,358.7

08:11 IE=+20 IA=-150: 362.5,358.5

But we don't see the star on IRTC........

*08:14 spiral search, found star on IRTC at 08:19 (see pointing log)

08:20 Turbulence on pupil seems to be going in a spiral pattern.?!
08:30 Open left side vent door.

08:42 SX Primary global offsets at EL=75 X=-1.65 Y=-2.9 Z=0.4 RX=0 RY=0

09:03 Z11 up to -1400 (net -400) we've been adjusting spherical up and down a bit in various pairs. Images in H are subarcsec! (seeing monitor is reporting 1 arcsec.)

*09:25 Pointing data in 022502 at EL=82 degrees SX primary global offsets X=-1.65 Y=-2.9 Z=+0.63 RX=0 RY=0

We seem to get this error from irs each time we take a picture after exiting continuous frame mode. irc GetIRParam > Problem with GetIRParam command. res=1

*09:49 Pointing data in 024923 at EL=84.1 degrees (some astigmatism)

09:55 Tested optical axis of IRTC by shifting between wide and medium scales. It looks quite good, how good is limited by the open-loop tracking drift (approx 0.1 arcsec/sec).

10:00 pointto WT10_378 16:35:41.302 57:46:51.76 but telescope went to 15:57 +26:52 which was our previous target. ??? irs reported no errors. And said:
[I] PresetTelescope command status: SUCCESS
[I] PCS set new target to RA: 4.178260633409109e+00 Dec: 4.691047966955929e-01 complete (That is DEC of +26.8)

10:04 repeated my pointo and PCS did a slewtotrack, but still PCS has the old target.

10:07 A preset from the presetGUI worked to 16:00 +35

10:09 pointto WT10_378 16:35:41.302 57:46:51.76 but telescope went BACK to 15:57 +26:52 AGAIN.

10:13 pointto WT10_266 16:54:10.066 15:37:24.33

target.dat file is missing the decimal point
Manual edit of the target.dat file allows the telescope to preset.

Called Jose to try to understand this. A failure in the SetStars command will leave the old target in place. But the ClearStars call at the beginning of Dave's pointto script should have emptied the old target.

Action to Jose: Understand the ClearStars issue (He figured out that he was checking guide stars flag but not target flag.) We'll get a new build of irs tomorrow.

Action to Dave: Put in the wavelength decimal point for pointto.cl

10:40 Verified that the PCS craziness when you send an IIF offset is proportional to the offset vector, and goes away if you zero out the relative offset.

Preset to WT10_266 via the preset GUI

11:02 No polys for 1 sec, telescope stops tracking in all three axes.

11:05 Telescope stops tracking again

11:15 Stop and restart PCS, MCS.

11:19 Slew Algorithm looks like it made a 17 deg error on the way back to our target.

Reloaded the pointing model.

*11:27 well-centered WT_266 at 72.6 deg EL in 042747 SX primary global offsets X=-1.6 Y=-3.3 Z=0.55 RX=0 RY=0

11:31 Preset to WT10_0411 (Note that presets from the preset GUI don't collimate automatically, so they aren't anymore real presets.)

*11:53 well-centered WT10_0411 at EL=38.1 roughly collimated in poor seeing. SX primary global offsets X=-2.1 Y=-3.3 Z=0.2 RX=0 RY=0

But maybe pointing log file wasn't running after we restarted PCS. Started it again "LAO"

*11:57 repeated the 11:53 image when we are sure pointing log was running.

11:58 Took a bunch of 0.5 sec dark exposures

12:03 Took some 0.05 sec dark sky exposures

12:10 Close Dome

Activities for Tomorrow
  • Analyze collimation data
  • Analyze pointing data
  • Learn how to run (turn on/turn off) IRTC before Italo leaves
  • Collect more collimation data
  • Take intra/extra focal pairs on fainter stars, or average multiple frames.
  • Fire up the AGw and see if it is in decent shape for the next steps.
  • Make the simple script for a spiral search (30 arcsec steps)

-- JohnHill - 13 Apr 2008
Topic revision: r8 - 16 Apr 2008, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback